FinanceCo lent $8.8 million to Corbin Construction on January 1, 2021, to construct a playground. Corbin signed a three-year, 5% installment note to be paid in three equal payments at the end of each year. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1)

Required:

1. Prepare the journal entry for FinanceCo’s lending the funds on January 1, 2021.

2. Prepare an amortization schedule for the three-year term of the installment note.

3. Prepare the journal entry for the first installment payment on December 31, 2021.

4. Prepare the journal entry for the third installment payment on December 31, 2023.
Answer #1

1)

Date Account title Debit credit
1 Jan 2021 Note receivable 8,800,000
cash 8,800,000

2)Annual installment = Amount provided /PVA5%,3

                              = 8800000 /2.72325

                              = 3231433 Rounded

Date Installment payment Interest expense principal repayment carrying value of note
1 Jan 2021 8800000
31 Dec 2021 3231433 8800000*5%= 440000 3231433-440000= 2791433 8800000-2791433= 6008567
31 Dec 2022 3231433 6008567*5%= 300428 3231433-300428= 2931005 6008567-2931005= 3077562
31 Dec 2023 3231433 3077562*5%= 153871 * 3077562 0

*Adjust for rounding of decimal places such the carrying value of note at end of year 3 is 0

3 and 4)In books of FinanceCo

Date Account title Debit credit
December 31 2021 Cash 3231433
Note receivable 2791433
interest revenue 440000
December 31 2023 Cash 3231433
Note receivable 3077562
interest revenue 153871